OpenStudy (konradzuse):

Integral question when you have something minus something else i.e., integral 48-12x^2.

OpenStudy (konradzuse):

Should you always break it up into \[\int\limits - \int\limits\] \[\int\limits 48 dx - \int\limits 12x^2 dx\]

OpenStudy (anonymous):

Usually makes it easier.

OpenStudy (anonymous):

no you don't have to, people do this so it becomes easier for them to see what to do, but it's not a law you have to apply.

OpenStudy (anonymous):

Can't think of a counterexample, though.

OpenStudy (konradzuse):

Okay, because the question was find the average of that from -5<x<5

OpenStudy (konradzuse):

so it became \[\frac{1}{10} \int\limits 48-12x^2dx\]

OpenStudy (konradzuse):

I figured I could take a 12 out so it would become \[\frac{6}{5} \int\limits 4-x^2dx\]

OpenStudy (konradzuse):

u = 4-x^2dx 1/2 du = xdx

OpenStudy (anonymous):

why would you use a substitution here?

OpenStudy (anonymous):

Hold on a second

OpenStudy (konradzuse):

Actually, I guess I could just split them up again huh.

OpenStudy (konradzuse):

\[\int\limits 4 - \int\limits x^2\]

OpenStudy (anonymous):

also don't forget the domains of your integral, you want to evaluate it.

OpenStudy (konradzuse):

-5 to 5

OpenStudy (konradzuse):

Ic, so that's where I made my mistake I guess.

OpenStudy (konradzuse):

so 4x-x^3/3 from -5 to 5

OpenStudy (anonymous):

So @KonradZuse slow down a second OK?

OpenStudy (konradzuse):

sure.

OpenStudy (anonymous):

Sorry I'm slow with the equation editor.

OpenStudy (anonymous):

You have the integral of 48 - 12x^2 between limits -5 to 5, right?

OpenStudy (anonymous):

Why not leave the constants in for now. The best thing to do is to take the two terms separately as you did. integral of 48 dx - integral of 12 x^2 dx right?

OpenStudy (konradzuse):

yup.. I know what the answer to that is, I was just trying to figure out if we got the same answer goign the other way is all...

OpenStudy (anonymous):

What would be the other way? How would you integrate (48 + 12x^2) dx? You don't have x

OpenStudy (anonymous):

Sorry, maybe I just got in your way. Anyhow, I haven't thought of a situation where you wouldn't split em (unless somehow factoring gave you something you needed for a substitution ..

OpenStudy (anonymous):

maybe I am now confused myself, it's getting late here, but integrating (48+12x^2)dx directly is possible

OpenStudy (anonymous):

Um.. let u = 48 + 12^2 You don't have the x for du

OpenStudy (anonymous):

well I understand that, so maybe we do that here a bit different then, but we let such integrals stay the way they are, keeping mind that the entire term in the brackets has to be integrated with respect to dx and then you just get the regular solution.

OpenStudy (anonymous):

I agree that it's not mathematically rigorous, but safes at least some writing for small terms.

OpenStudy (anonymous):

Maybe I don't understand your example. Can you show me?

OpenStudy (anonymous):

\[ \int (48 +12x^3)dx\] Brackets just say that everything inside is multiplied by dx, but instead of writing that out you can just do it directly, it's the same thing as you do, just wont require splitting it up, but as I said, it loses some rigor definitions about the integral and its function as a linear operator. \[ \int 48dx+12x^3dx= 48x+3x^4\]

OpenStudy (anonymous):

Well, OK. You are splitting it up, just saving yourself actually writing the second integral sign. I can live with that. Personally I find the parentheses confusing because then I'm thinking that is u(x) and I need du..

OpenStudy (anonymous):

Yes I understand that, I see the above method applied all the time, without even carrying out the dx (second step I did), but yes it is mainly for saving some time for 'easier' integrals like this, linear exponent terms. I just asked because it's interesting for me how people apply various rules over the globe (-: there are some little differences.

OpenStudy (anonymous):

Great chatting. Get some sleep.

OpenStudy (konradzuse):

Thanks goodniught.